OpenStudy (anonymous):

Write this in slope intercept form A slope of 0 and containing the point (-1,7) Y=_________?

OpenStudy (precal):

y=7

OpenStudy (anonymous):

Yes, since (x,y); y=7

OpenStudy (anonymous):

7 was not corret

OpenStudy (anonymous):

@cwhutche are you looking for b, the y-interecept? Or the y=mx+b eqn??

OpenStudy (anonymous):

y=mxtb

OpenStudy (anonymous):

y - 7 = 0(x +1 y = 0 +7 y = 7 y = 0x + 7

Directrix (directrix):

With a slope of 0, the line is horizontal. The point has y-coordinate 7. I'd say y = 7.

OpenStudy (anonymous):

i tried 7 but doesnt work

OpenStudy (anonymous):

it wont work

OpenStudy (precal):

 yes it does work, perhaps your online course is not accepting the solution because of how you input the solution